# Temporal reasoning for timeline summarisation in social media

### v2

@@ Abstract
This paper explores whether enhancing temporal reasoning capabilities in Large Language Models (LLMs) can improve the quality of timeline summarisation, the task of summarising long texts containing sequences of events, such as social media threads. We first introduce NarrativeReason, a novel dataset focused on temporal relationships among sequential events within narratives, distinguishing it from existing temporal reasoning datasets that primarily address pair-wise event relationships. Our approach then combines temporal reasoning with timeline summarisation through a knowledge distillation framework, where we first fine-tune a teacher model on temporal reasoning tasks and then distill this knowledge into a student model while simultaneously training it for the task of timeline summarisation. Experimental results demonstrate that our model achieves superior performance on out-of-domain mental health-related timeline summarisation tasks, which involve long social media threads with repetitions of events and a mix of emotions, highlighting the importance and generalisability of leveraging temporal reasoning to improve timeline summarisation.
@@ 1 Introduction
Temporal reasoning involves understanding and processing temporal information in text to deduce time-based relationships between events Wenzel and Jatowt (2023a). Zhou et al. (2019) categorises temporal commonsense reasoning with respect to five aspects (duration, temporal ordering, typical time, frequency and stationarity). Subsequently, Tan et al. (2023); Jain et al. (2023) explore the temporal reasoning capabilities of Large Language Models (LLMs) with respect to temporal commonsense aspects. LLMs with a strong understanding of temporal context can perform better on downstream tasks, including storytelling, natural language inference, timeline comprehension and tracking user status Jain et al. (2023). Thus temporal commonsense reasoning is beneficial for timeline summarisation, as it helps maintain temporal consistency and the correct event order Wenzel and Jatowt (2023b); Vashishtha et al. (2020). Despite the evidenced connection between temporal reasoning and timeline summarisation, recent work Chan et al. (2024); Feng et al. (2023); Chen et al. (2024); Zhang et al. (2024) has primarily focused on improving the temporal reasoning capabilities of LLMs, without exploring how they impact downstream tasks, such as timeline summarisation.
@@
    Based on the timelines derived in NarrativeTime Rogers et al. (2024) from the TimeBankNT corpus Cassidy et al. (2014), we develop a new dataset for temporal reasoning, NarrativeReason. Unlike existing temporal reasoning datasets Tan et al. (2023); Chu et al. (2024); Wang and Zhao (2024), NarrativeReason focuses on the temporal relationships among a series of events within a story rather than distinct event pairs. This can help LLMs process a series of events to generate a coherent and accurate timeline summary.
@@
    We fine-tune a large LLM on NarrativeReason and distill its knowledge to a smaller model, which is fine-tuned for the task of timeline summarisation. The resulting fine-tuned smaller LLM is applied to a completely different domain from the one it is trained on. Experimental results show that our model achieves the best performance on the timeline summarisation dataset by Song et al. (2024). Not only does it generate more accurate summaries but it also reduces hallucinations in LLMs.
@@ 2 Related Work
Temporal reasoning for LLMs Temporal reasoning in Natural Language Processing (NLP) is the ability to understand and process information related to time within natural language text. It includes reasoning about the chronology and duration of events, and understanding and capturing different temporal relations Vashishtha et al. (2020). Despite the impressive performance of Large Language Models (LLMs), like GPT-4, across a wide range of tasks (e.g. translation, generation), they have been shown to perform sub-optimally in temporal reasoning Wang and Zhao (2024); Chu et al. (2024); Qiu et al. (2023). However the ability to perform temporal reasoning is crucial for understanding narratives Nakhimovsky (1987); Jung et al. (2011a); Cheng et al. (2013), answering questions Bruce (1972); Khashabi (2019); Ning et al. (2020), and summarising events Jung et al. (2011b); Vashishtha et al. (2020). Consequently, efforts are being made to enhance the temporal reasoning capabilities of LLMs Xing and Tsang (2023) Huang et al. (2024). To increase understanding of temporal expressions, Tan et al. (2023) introduced the TEMPREASON dataset which addresses three types of relations (time-time, time-event, event-event). TEMPREASON was used to fine-tune a LLM to improve its temporal reasoning, and performance of different LLMs on this dataset showed it is challenging for LLMs to capture the temporal relations between different events. Xiong et al. (2024) use an aligned timeline to improve an LLM’s temporal reasoning by translating the context into a temporal graph,identifying valid time expressions and generating related temporal knowledge. The temporal relationship between events is inferred based on specific times (e.g., year of event). However in a narrative, events often occur without a clear indication of time.
Temporal reasoning for summarisation Jung et al. (2011b) developed a natural language understanding (NLU) system with a temporal reasoning component to create comprehensive timelines, applied to medical records, presenting medical history in a more intuitive way. They found that temporal reasoning in NLU is tightly integrated into the NLP system’s deep semantic analysis and can help a LM analyze temporal relationships between different events, which is beneficial for event or news summarisation Vashishtha et al. (2020). However, despite the evidenced connection, few studies explore how improvements in temporal reasoning in LLMs directly benefit downstream tasks such as text summarisation.
@@ 3.2 Teacher Model and NarrativeReason dataset
Event extraction To create NarrativeReason we restructured the NarrativeTime dataset Rogers et al. (2024), which in turn had re-annotated TimeBankDense Cassidy et al. (2014) with a timeline-based annotation framework, NarrativeTime framework. Rogers et al. (2024) had annotated all possible temporal links (TLINKS) between all events occurring within a narrative, thus providing temporal relations for the entire sequence of events (timeline) rather than just between event pairs to get a NarrativeTime dataset. Here, TLINKs usually denote the event order information (e.g., before, after, during). Events more broadly and especially in the context of temporal reasoning, are represented as relation triples where the event trigger is the head of the verb phrase Ning et al. (2018); Pustejovsky et al. (2003), linking the corresponding arguments. However, NarrativeTime dataset  Rogers et al. (2024) only denotes the type of the event, annotated at the level of the verb, e.g. for the sentence "the value of the Indonesian stock market has fallen by twelve percent" this is annotated in NarrativeTime datasets as "the value of the Indonesian stock market has <EVENT class="OCCURRENCE" eid="e7"> fallen </EVENT> by twelve percent". However, this only constitutes a denser temporal annotation without providing event triples and is therefore unsuited for temporal reasoning training tasks.
Thus we reconstructed NarrativeReason, to augment it with event triples that can be used for temporal reasoning training. Specifically, we filtered the NarrativeTime dataset, keeping only annotated verbs to represent events. In order to represent an event, we use verbs as triggers to extract relational triples e.g. <Indonesian stock market value, fallen, by twelve percent> to represent the event annotated in NarrativeTime dataset as fallen. Then, we use these triples to construct the temporal relations of a series event. (e.g. Event <Indonesian stock market value, fallen, by twelve percent> is BEFORE Event <financial week, turning, bad for Asia>).
@@
Fine-tuning task We apply supervised fine-tuning (SFT) on a large LLM (teacher model) utilising Low-Rank Adaptation (LoRA) Hu et al. (2022). The input and output of the model are the temporal questions and corresponding answers. Our experiments show that indeed fine-tuning on NarrativeReason improves performance on established temporal reasoning tasks (Appendix A.2) .
Figure 2: The temporal relations between events. The text in the left column comes from NarrativeTime dataset, and we filtered it to keep only verbs. We represent events triggered by verbs using relational triples, as shown in the middle column. In the right column, we construct question/answer pairs for all events.
@@ 3.3 Student Model
where ${K(x,y)=exp(-\frac{{||x-y||}_{2}^{2}}{2{\sigma}^{2}})}$ with $\sigma=1$ is Gaussian Kernel. We transfer the teacher activation patterns to the student by minimizing $L_{MMD^{2}}$.
@@
The knowledge distillation process transfers temporal reasoning knowledge from the teacher to the student model. At the same time, we want this knowledge to benefit the timeline summarisation task. Thus, we fine-tune the student model on the timeline summarisation dataset §4.1, enabling it to both learn from the teacher model and use the language modeling loss (for next token prediction) $L_{language}$ to integrate temporal reasoning knowledge with timeline summarisation information.
@@ 3.4 Mental Health Timeline Summary
We apply the student model to other domains, specifically to generate mental health-related summaries for timelines §4.1 from social media. For the mental health summary, we use the format proposed by Song et al. (2024), which includes three key clinical concepts (diagnosis, inter- and intra- personal relations, moments of change). We follow their method to prompt the student model to generate a summary for each timeline.
@@ 4.1 Datasets
NarrativeReason We extracted 668 events from 30 articles, containing a total of 19,614 temporal relations between events Rogers et al. (2024), leading to 19,614 question/answer pairs for event-event relations. We use these question/answer pairs to fine-tune the teacher model to enhance its temporal reasoning capability.
@@ 4.2 Models & Baselines
P-Phi: The same as L-Phi but we use another Phi as the teacher model instead of LLaMA-3.
@@
TH-VAE: We use TH-VAE Song et al. (2024) to generate mental health summaries for comparison, while we use LLaMA-3 to generate high-level summaries.
@@ 4.3 Evaluation
We work with the timeline summaries from Song et al. (2024)4.1). Like Song et al. (2024), we employ Factual Consistency (FC), to measure how consistent timeline summaries are with the original timelines, and Evidence Appropriateness (EA), to measure the consistency between human written summaries and corresponding timeline summaries Song et al. (2024). Here, we use the annotated timeline evidence from Song et al. (2024) to directly generate high-level summaries. By contrast Song et al. (2024), generated clinically meaningful summaries from the timeline by first using TH-VAE to generate evidence related to mental health before generating high-level summaries. Since the goal of our paper is to validate the role of temporal reasoning in timeline summarisation, we directly use the annotated evidence to generate the high-level summary as the point is to make better use of the evidence rather than extract it from scratch.  [^2] We obtained the annotated dataset from the authors Song et al. (2024).
@@ 5 Results and Discussion
Automatic evaluation: We conducted experiments with different combinations of KD methods and fine-tuning strategies (Table 1). Since we didn’t change output dimensions when fine-tuning LLaMA, CRD was not used during the KD process. Among individual methods, PKT performed the best, but combining PKT with NST achieved the best overall results. P-Phi and L-Phi are the best performing models with identical EA (correspondence to human summaries). L-Phi seems to have lower FC, which would suggest it is less faithful to the original timeline. However, human evaluation in Table 2 indicates that clinical psychologists have a clear preference for summaries generated by L-Phi, on all aspects. This suggests distilling information from a larger LLM is indeed beneficial.
@@
Table 1: Automatic evaluation for factual consistency (FC), evidence appropriateness (EA) for timeline summarisation of the different models. Higher is better. Best results are in bold and second-best results in underline.
Regarding fine tuning strategies for timeline summarisation, the results for $\text{Phi}_{temp}$ and $\text{Phi}_{tl}$ on FC indicate that fine-tuning on a single dataset does not improve model performance; instead, it exacerbates hallucination issues. Notably, $\text{Phi}_{temp}$ performed the worst on both FC and EA metrics, suggesting that incorporating temporal reasoning information interferes with the LLM’s ability to effectively handle the timeline summarisation task. Additionally, in $\text{Phi}_{joint}$, combining the two types of data directly during training failed to integrate them effectively. As a result, the performance of $\text{Phi}_{joint}$ was worse than just using in-context learning to guide the LLM ($\text{Phi}_{ICL}$).
Human evaluation: Based on the results of the automatic evaluation, we selected the best-performing L-Phi and P-Phi models. Additionally, we included the non-fine-tuned zero-shot versions of LLaMA and Phi. This can help us understand in which specific aspects the model has improved with the inclusion of temporal reasoning information. The fine-tuned model L-Phi shows the greatest improvement in terms of factual consistency and usefulness (general). This aligns with our findings when analyzing the summaries, where the fine-tuned model significantly reduces hallucination, as shown in Table 2. In addition, we found that the fine-tuned model did not show significant improvement in terms of Moments of Change (MOC). In Appendix  A.4, we give examples of summaries generated by different models, these examples clearly demonstrate that the fine-tuned model can effectively reduce hallucinations.
@@ 5.1 Why knowledge distillation works
To calculate CKA, we used the same probing dataset. First, we calculate the sentence embeddings of each input by averaging the hidden state representation of the tokens. Then, we calculate the CKA similarity score between the mean sentence embeddings and each layer representation of the model. We did this both for the individual models and between models’ layers. From Figure 4, we can see that $\text{Phi}_{tl}$ shows a gradual increase in CKA across layers, peaking in the mid-to-late layers. This indicates that as the layers progress, $\text{Phi}_{tl}$ preserves and refines the information from the initial embeddings in a task-relevant way. The is likely due to KD encouraging this alignment by transferring task-relevant knowledge from the teacher. While $\text{Phi}_{joint}$ shows an initial increase in CKA, it saturates and flattens early. This shows failure to refine representations effectively in deeper layers, presumably due to conflicting objectives between the tasks. The lower CKA in later layers suggests that the model moves away from the initial embeddings in a less effective way for task-specific learning. Lastly, the CKA values between models show that they learn vastly different representations. In short, $\text{Phi}_{tl}$’s ability to align with the initial embeddings correlates with its better task performance, as the CKA value reflects how well the model retains and transforms meaningful input information throughout its layers.
(a) UMAP projection $\text{Phi}_{joint}$
(b) UMAP projection $\text{Phi}_{tl}$
[…]
Figure 4: CKA similarity score of both within and between $\text{Phi}_{tl}$ and $\text{Phi}_{joint}$ model representations.

### v3

@@ Abstract
This paper explores whether enhancing temporal reasoning capabilities in Large Language Models (LLMs) can improve the quality of timeline summarisation, the task of summarising long texts containing sequences of events, such as social media threads. We first introduce NarrativeReason, a novel dataset focused on temporal relations among sequential events within narratives, distinguishing it from existing temporal reasoning datasets that primarily address pair-wise event relations. Our approach then combines temporal reasoning with timeline summarisation through a knowledge distillation framework, where we first fine-tune a teacher model on temporal reasoning tasks and then distill this knowledge into a student model while simultaneously training it for the task of timeline summarisation. Experimental results demonstrate that our model achieves superior performance on out-of-domain mental health-related timeline summarisation tasks, which involve long social media threads with repetitions of events and a mix of emotions, highlighting the importance and generalisability of leveraging temporal reasoning to improve timeline summarisation.
@@ 1 Introduction
Temporal reasoning involves understanding and processing temporal information in text to deduce time-based relations between events Wenzel and Jatowt (2023a). Zhou et al. (2019) categorises temporal commonsense reasoning with respect to five aspects (duration, temporal ordering, typical time, frequency and stationarity). Subsequently, Tan et al. (2023); Jain et al. (2023) explore the temporal reasoning capabilities of Large Language Models (LLMs) with respect to temporal commonsense aspects. LLMs with a strong understanding of temporal context can perform better on downstream tasks, including storytelling, natural language inference, timeline comprehension and tracking user status Jain et al. (2023). Thus temporal commonsense reasoning is beneficial for timeline summarisation, as it helps maintain temporal consistency and the correct event order Wenzel and Jatowt (2023b); Vashishtha et al. (2020). Despite the evidenced connection between temporal reasoning and timeline summarisation, recent work Chan et al. (2024); Feng et al. (2023); Chen et al. (2024); Zhang et al. (2024) has primarily focused on improving the temporal reasoning capabilities of LLMs, without exploring how they impact downstream tasks, such as timeline summarisation.
@@
    Based on the timelines derived in NarrativeTime Rogers et al. (2024) from the TimeBankNT corpus Cassidy et al. (2014), we develop a new dataset for temporal reasoning, NarrativeReason. Unlike existing temporal reasoning datasets Tan et al. (2023); Chu et al. (2024); Wang and Zhao (2024), NarrativeReason focuses on the temporal relations among a series of events within a story rather than distinct event pairs. This can help LLMs process a series of events to generate a coherent and accurate timeline summary.
@@
    We fine-tune a large LLM on NarrativeReason and distill its knowledge to a smaller model, which is fine-tuned for the task of timeline summarisation. The resulting fine-tuned smaller LLM is applied to a completely different domain from the one it is trained on. Experimental results show that our model achieves the best performance on the timeline summarisation dataset by Song et al. (2024). Not only does it generate more accurate summaries, but it also reduces hallucinations in LLMs.
@@ 2 Related Work
Temporal reasoning for LLMs Temporal reasoning in Natural Language Processing (NLP) is the ability to understand and process information related to time within natural language text. It includes reasoning about the chronology and duration of events, and understanding and capturing different temporal relations Vashishtha et al. (2020). Despite the impressive performance of Large Language Models (LLMs), like GPT-4, across a wide range of tasks (e.g. translation, generation), they have been shown to perform sub-optimally in temporal reasoning Wang and Zhao (2024); Chu et al. (2024); Qiu et al. (2023). However the ability to perform temporal reasoning is crucial for understanding narratives Nakhimovsky (1987); Jung et al. (2011a); Cheng et al. (2013), answering questions Bruce (1972); Khashabi (2019); Ning et al. (2020), and summarising events Jung et al. (2011b); Vashishtha et al. (2020). Consequently, efforts are being made to enhance the temporal reasoning capabilities of LLMs Xing and Tsang (2023) Huang et al. (2024). To increase understanding of temporal expressions, Tan et al. (2023) introduced the TEMPREASON dataset which addresses three types of relations (time-time, time-event, event-event). TEMPREASON was used to fine-tune a LLM to improve its temporal reasoning, and performance of different LLMs on this dataset showed it is challenging for LLMs to capture the temporal relations between different events. Xiong et al. (2024) use an aligned timeline to improve an LLM’s temporal reasoning by translating the context into a temporal graph,identifying valid time expressions and generating related temporal knowledge. The temporal relations between events are inferred based on specific times (e.g., year of event). However in a narrative, events often occur without a clear indication of time.
Temporal reasoning for summarisation Jung et al. (2011b) developed a natural language understanding (NLU) system with a temporal reasoning component to create comprehensive timelines, applied to medical records, presenting medical history in a more intuitive way. They found that temporal reasoning in NLU is tightly integrated into the NLP system’s deep semantic analysis and can help a LM analyze temporal relations between different events, which is beneficial for event or news summarisation Vashishtha et al. (2020). However, despite the evidenced connection, few studies explore how improvements in temporal reasoning in LLMs directly benefit downstream tasks such as text summarisation.
@@ 3.2 Teacher Model and NarrativeReason dataset
Event extraction To create NarrativeReason we restructured the NarrativeTime dataset Rogers et al. (2024), which in turn had re-annotated TimeBankDense Cassidy et al. (2014) with a timeline-based annotation framework, NarrativeTime. Rogers et al. (2024) have annotated all possible temporal links (TLINKS) between all events occurring within a narrative, thus providing temporal relations for the entire sequence of events (timeline) rather than just between event pairs to get obtain the NarrativeTime dataset. Here, TLINKs usually denote the event order information (e.g., before, after, during). Events more broadly and especially in the context of temporal reasoning, are represented as relation triples where the event trigger is the head of the verb phrase Ning et al. (2018); Pustejovsky et al. (2003), linking the corresponding arguments. However, NarrativeTime  Rogers et al. (2024) only denotes the type of event, annotated at the level of the verb, e.g. for the sentence "the value of the Indonesian stock market has fallen by twelve percent" this is annotated in the NarrativeTime dataset as "the value of the Indonesian stock market has <EVENT class="OCCURRENCE" eid="e7"> fallen </EVENT> by twelve percent". However, this only constitutes a denser temporal annotation without providing event triples and is therefore unsuited for temporal reasoning training tasks.
Thus we reconstructed NarrativeReason, to augment it with event triples that can be used for temporal reasoning training. Specifically, we filtered the NarrativeTime dataset, keeping only annotated verbs to represent events. In order to represent an event, we use verbs as triggers to extract relational triples e.g. <Indonesian stock market value, fallen, by twelve percent> now represents the event annotated in the NarrativeTime dataset as fallen. Then, we use these triples to construct the temporal relations of a series of events. (e.g. Event <Indonesian stock market value, fallen, by twelve percent> is BEFORE Event <financial week, turning, bad for Asia>).
@@
Fine-tuning task We apply supervised fine-tuning (SFT) on a large LLM (teacher model) utilising Low-Rank Adaptation (LoRA) Hu et al. (2022). The input and output of the model are the temporal questions and corresponding answers respectively. Our experiments show that indeed fine-tuning on NarrativeReason improves performance on established temporal reasoning tasks (Appendix A.2) .
Figure 2: The temporal relations between events. The text in the left column comes from the NarrativeTime dataset, and we filtered it to keep only verbs. We represent events triggered by verbs using relational triples, as shown in the middle column. In the right column, we construct question/answer pairs for all events.
@@ 3.3 Student Model
where ${K(x,y)=exp(-\frac{{||x-y||}_{2}^{2}}{2{\sigma}^{2}})}$ with $\sigma=1$ is the Gaussian Kernel. We transfer the teacher activation patterns to the student by minimizing $L_{MMD^{2}}$.
@@
The knowledge distillation process transfers temporal reasoning knowledge from the teacher to the student model. At the same time, we want this knowledge to benefit the timeline summarisation task. Thus, we fine-tune the student model on the timeline summarisation dataset §4.1, enabling it to both learn from the teacher model and use the language modeling loss $L_{language}$ (for next token prediction) to integrate temporal reasoning knowledge with timeline summarisation information.
@@ 3.4 Mental Health Timeline Summary
We apply the student model to other domains, specifically to generate mental health-related summaries for timelines §4.1 from social media. For mental health summaries, we use the format proposed by Song et al. (2024), which includes three key clinical concepts (diagnosis, inter- and intra- personal relations, moments of change). We follow their method to prompt the student model to generate a summary for each timeline.
@@ 4.1 Datasets
NarrativeReason We extracted 668 events from 30 articles, containing a total of 19,614 temporal relations between events in Rogers et al. (2024), leading to 19,614 question/answer pairs for event-event relations. We use these question/answer pairs to fine-tune the teacher model to enhance its temporal reasoning capability.
@@ 4.2 Models & Baselines
P-Phi: The same configuration as L-Phi but we use another Phi as the teacher model instead of LLaMA-3.
@@
$\text{KD}_{timeline}$: To assess whether temporal reasoning knowledge contributes to the observed positive improvements, we fine-tune the teacher with the timeline summarisation dataset, and transfer it to the student model.
$\text{KD}_{origin}$: We also transfer knowledge from the teacher model without finetuning on any dataset. Instead, we examine whether the teacher’s inherent knowledge can help improve the timeline summarization task.
@@
TH-VAE: For comparison with the state-of-the-art on this dataset we use TH-VAE Song et al. (2024) to generate mental health timeline summaries which are then translated by LLaMA-3 to high-level summaries.
@@ 4.3 Evaluation
We work with the timeline summaries from Song et al. (2024)4.1). Like Song et al. (2024), we employ Factual Consistency (FC), to measure how consistent timeline summaries are with the original timelines, and Evidence Appropriateness (EA), to measure the consistency between human written summaries and corresponding timeline summaries Song et al. (2024). FC and EA combined effectively capture whether the generated summary aligns with factual information. Thus increase in these scores means reduction in hallucinations and generation of more reliable summaries. Here, we use the annotated timeline evidence from Song et al. (2024) to directly generate high-level summaries. By contrast Song et al. (2024), generated clinically meaningful summaries from the timeline by first using TH-VAE to generate evidence related to mental health before generating high-level summaries. Since the goal of our paper is to validate the role of temporal reasoning in timeline summarisation, we directly use the annotated evidence in the dataset to generate the high-level summary as the point is to make better use of the evidence rather than extract it from scratch.
@@ 5 Results and Discussion
Automatic evaluation: We conducted experiments with different combinations of KD methods (Table 1). Since we didn’t change output dimensions when fine-tuning LLaMA, CRD was not used during the KD process. Among individual methods, PKT performed the best, but combining PKT with NST achieved the best overall results. We know that NST matches the distributions of neuron selectivity patterns between teacher and student networks; PRT also matches the probability distribution of the data in the feature space of teacher and student models. However, CRD focuses on distinguishing between different instances rather than learning structural consistency. In learning temporal reasoning, contrastive learning (CRD) may not be as effective as NST and PRT, since the task requires maintaining temporal coherence rather than separating instances.
P-Phi and L-Phi are the best performing models with identical EA (correspondence to human summaries). L-Phi seems to have lower FC, which would suggest it is less faithful to the original timeline. However, human evaluation in Table 3 indicates that clinical psychologists have a clear preference for summaries generated by L-Phi, on all aspects. This suggests distilling information from a larger LLM is indeed beneficial.
@@
Table 1: Automatic evaluation for factual consistency (FC), evidence appropriateness (EA) for timeline summarisation of the different KD strategies. Higher is better. Best results are in bold and second-best results in underline.
[…]
Table 2: Automatic evaluation for factual consistency (FC), evidence appropriateness (EA) for timeline summarisation of all the different models we consider. Higher is better. Best results are in bold and second-best results in underline.
Regarding fine tuning strategies for timeline summarisation as shown in Table 2 , the results for $\text{Phi}_{temp}$ and $\text{Phi}_{tl}$ on FC indicate that fine-tuning on a single dataset does not improve model performance; instead, it exacerbates hallucination issues. Notably, $\text{Phi}_{temp}$ performed the worst on both FC and EA metrics, suggesting that incorporating temporal reasoning information interferes with the LLM’s ability to effectively handle the timeline summarisation task. Additionally, in $\text{Phi}_{joint}$, combining the two types of data directly during training, failed to integrate them effectively. As a result, the performance of $\text{Phi}_{joint}$ was worse than just using in-context learning to guide the LLM ($\text{Phi}_{ICL}$).
[…]
Human evaluation: Based on the results of the automatic evaluation, we selected the best-performing L-Phi and P-Phi models. Additionally, we included the non-fine-tuned zero-shot versions of LLaMA and Phi. This can help us understand in which specific aspects the model has improved with the inclusion of temporal reasoning information. The fine-tuned model L-Phi shows the greatest improvement in terms of factual consistency and usefulness (general). This aligns with our findings when analyzing the summaries, where the fine-tuned model significantly reduces hallucination, as shown in Table 3. In addition, we found that the fine-tuned model did not show significant improvement in terms of Moments of Change (MOC). In Appendix  A.4, we give examples of summaries generated by different models. These examples clearly demonstrate that the fine-tuned model can effectively reduce hallucinations.
@@ 5.1 Why knowledge distillation works
Here we analyze from a representation learning perspective why the L-Phi model performs better. We run two experiments on: (1)task understanding probing experiment and (2) Joint Task Representation Learning (JTRL). We analyze the internal representations of L-Phi against $\text{Phi}_{joint}$. We construct our probing dataset from the test dataset of the latter. We pose the two tasks as binary classification and extract activations for the last layer. We use UMAP (McInnes et al., 2018) to project the activations to lower dimensions (Sainburg et al., 2021; Tseriotou et al., 2023). In Figure 3, we can see that the activations of $\text{Phi}_{joint}$ are well separated for each task, whereas those of L-Phi overlap. Given the performance of L-Phi, this would suggest that the model learned better representations for the tasks due to more task-specific polysemantic (Olah et al., 2020) neurons.
To calculate CKA, we used the same probing dataset. First, we calculate the sentence embeddings of each input by averaging the hidden state representation of the tokens. Then, we calculate the CKA similarity score between the mean sentence embeddings and each layer representation of the model. We did this both for the individual models and between models’ layers. From Figure 4, we can see that L-Phi shows a gradual increase in CKA across layers, peaking in the mid-to-late layers. This indicates that as the layers progress, L-Phi preserves and refines the information from the initial embeddings in a task-relevant way. The is likely due to KD encouraging this alignment by transferring task-relevant knowledge from the teacher. While $\text{Phi}_{joint}$ shows an initial increase in CKA, it saturates and flattens early. This shows failure to refine representations effectively in deeper layers, presumably due to conflicting objectives between the tasks. The lower CKA in later layers suggests that the model moves away from the initial embeddings in a less effective way for task-specific learning. Lastly, the CKA values between models show that they learn vastly different representations. In short, L-Phi’s ability to align with the initial embeddings correlates with its better task performance, as the CKA value reflects how well the model retains and transforms meaningful input information throughout its layers.
(a) UMAP projection $\text{Phi}_{joint}$
(b) UMAP projection L-Phi
@@
Figure 4: CKA similarity score of both within and between L-Phi and $\text{Phi}_{joint}$ model representations.